OpenStudy (anonymous):

What is the solution 2/4x^2 = 1/16x?

OpenStudy (anonymous):

I did all of the other questions on my test, the two I asked were the only two I didn't understand. XD

OpenStudy (blurbendy):

you have: 2/4x^2 = 1/16x which is equal to 2/4x^2 - 1/16x = 0 the denominators aren't common, so what do we do when we don't have a common denominator?

OpenStudy (anonymous):

We multiply to make them the same, but I'm not exactly sure what to do first.

OpenStudy (blurbendy):

okay, so what's the least common denominator for x^2 and x?

OpenStudy (anonymous):

x^2?

OpenStudy (blurbendy):

yup, so how do we make x = x^2?

OpenStudy (anonymous):

We multiply it by x, or square it?

OpenStudy (blurbendy):

multiply it x x*x = x^2 okay, so multiply 1/16x by x / x

OpenStudy (anonymous):

x/16x^2

OpenStudy (blurbendy):

now we have 2/4x^2 - x/16x^2 denominators still aren't common, so we need 4x^2 to equal 16x^2. how do we do that?

OpenStudy (anonymous):

Can we take 2/4x^2 and do the whole thing multiplied by 4?

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

So, 8/16x^2 = x/16x^2?

OpenStudy (blurbendy):

yup

OpenStudy (blurbendy):

you can simplify those a bit

OpenStudy (anonymous):

And so is the solution, when simplified, just 8?

OpenStudy (blurbendy):

you should get two solutions since it is x^2 i got 0 and 1/8

OpenStudy (blurbendy):

if you want to check, you can substitute 1/8 into the original equation, and you should get a true statement.

OpenStudy (anonymous):

So disregard the 0 since the 1/8 works?

OpenStudy (blurbendy):

0 should work too, but i dont see that as an answer choice

OpenStudy (anonymous):

I just turned that test in and it said 1/8 was wrong... :O

OpenStudy (anonymous):

@blurbendy

OpenStudy (blurbendy):

impossibleeee!

OpenStudy (anonymous):

So was the first question I asked... Now I'm really confused...

OpenStudy (blurbendy):

if you put 1/8 into the original equation, you get an equality. i have no idea why that was wrong

OpenStudy (blurbendy):

did you forget a negative sign or something in the original equation?

OpenStudy (anonymous):

I can reset my exam and take it again... Hopefully I'll do better the second time around. XD

OpenStudy (blurbendy):

that doesnt make any sense. im mad now

OpenStudy (anonymous):

Me too!

OpenStudy (blurbendy):

can you send this to your teacher or something? the computer is wrong!

OpenStudy (blurbendy):

send him/her a picture and say that if you substitute 1/8 into x, then you get a true statement...

OpenStudy (anonymous):

Yeah, I'll email her. And the first question was correct, too, right? This is really... ridiculous.

OpenStudy (blurbendy):

you got that one wrong too? wow.... that's some faulty software.

OpenStudy (anonymous):

Yes! And this was my last assignment before finishing the class. I have an A, but I really wanted to go out with a bang. .__.

OpenStudy (blurbendy):

haha nice. well good luck! let me know how it goes

OpenStudy (anonymous):

I will! Thanks for your help, though! :D